# Exploring Automated Distractor Generation for Math Multiple-choice Questions via Large Language Models

### v2

@@ Abstract
[…]
@@ 1.1 Contributions
  • We explore a variety of approaches to this task, including in-context learning, fine-tuning, and chain-of-thought prompting, together with rule- and sampling-based baselines.
@@ 3.3 Experimental Setup
For all approaches except SB, we use a uniform format to represent the target MCQ. This format comprises a concatenation of 3 elements: the question stem, key, and its explanation. We use this structure since it encapsulates the most comprehensive information about the target MCQ. Furthermore, based on CoT, we instruct the LLM to first generate feedback message and then the distractor, which intends to simulate a reasoning pathway, providing a scaffold that guides the subsequent generation of plausible distractors. We use greedy decoding and a maximum output length of 350 tokens for distractor generation. Additional hyperparameters and model details are in Supplementary Material Section B. We also provide our prompts for CoT, RB, and kNN in Tables  9,  10, and  11 respectively.
@@ 4 Conclusions and Future Work
In this paper, we explore automated distractor generation for math multiple-choice questions via large language models. We conduct experiments on a real-world math MCQ dataset and find that the in-context learning-based approach kNN, achieves the best performance when compared to other approaches such as fine-tuning, chain-of-thought prompting, and various baselines. We also conduct human evaluation and observe that LLMs are capable of generating mathematically valid distractors but are not fully aware of common errors or misconceptions among real students. Our initial exploration of this task opens up many avenues for future work. For example, we need to further refine the distribution-based metrics that predict the percentage of students who select each distractor. We also need to develop modified text encoding approaches that are closely aligned with errors or misconceptions among real students for in-context example selection. Furthermore, we aim to explore the generation of distractors, each of which corresponds to a specific error or misconception, as well as the generation of high-quality feedback messages for each distractor.
5 Acknowledgements
The authors thank Schmidt Futures and the NSF (under grants 2118706 & 2237676) for supporting this work.
@@ Appendix A Distribution ranking metric
Formally, we train an LLM-based model $\operatorname{r}_{\phi}(d_{1},d_{2},s,k,e_{k})\rightarrow\{d_{1},d_{2}\}$, where $\phi$ denotes the set of model parameters. We train this model by first constructing a dataset of all pairs of human-authored distractors for each MCQ and include both orders of each pair to avoid ordering bias, resulting in $N\times{3\choose 2}\times 2$ total pairs, where $N$ denotes the number of MCQs. Each pair is associated with a binary-valued label indicating whether $d_{1}$ or $d_{2}$ is selected by more students, which we can calculate from the student response records in our dataset. We then use this dataset to fine-tune an LLM in a text generation task, where the LLM receives the question and distractor information in its prompt and outputs its preference. We show our prompt for this task in Table 12.
@@ Appendix C Human Evaluation Details
In this work, we obtained approval from the ethics review board for human evaluation. We show the evaluation instructions to human evaluators in Table  13. We do not provide any compensation for human evaluators because their participation is entirely voluntary and we appreciate their contribution to this work.
@@ Appendix D Prompt Format
Prompt
You are given the following math question along with the correct answer and explanation. Please use the following template to give 3 alternative incorrect answers to be used as multiple-choice options in a multiple-choice exam. Prior to the incorrect answer, provide feedback to be displayed to the student as an explanation of why that is not the correct answer.
@@
Prompt
You are given the following math question along with the correct answer, explanation, and a list of errors. Please follow the template to first select 3 most likely errors for this question and use the selected errors to generate 3 alternative incorrect answers to be used as multiple-choice options in a multiple-choice exam. Prior to the incorrect answer, provide feedback to be displayed to the student as an explanation of why that is not the correct answer. If the list of errors is not given, generate 3 errors instead and do not contain any explanation in the 3 incorrect answer.
@@
Prompt
Question: <in-context question>
@@ Appendix E Ranking Metric Examples
Prompt
A teacher assigns the following math multiple choice question to a class of middle school students.
@@
Output
Preferred Answer: A

### v3

@@ Abstract
Wanyong Feng1∗, Jaewook Lee1∗, Hunter McNichols1∗, Alexander Scarlatos1∗, Digory Smith2, Simon Woodhead2, Nancy Otero Ornelas3, Andrew Lan1 University of Massachusetts Amherst1, Eedi2, Stanford University, Kitco Design3 {wanyongfeng,ajscarlatos,jaewooklee,wmcnichols,andrewlan}@umass.edu {digory.smith,simon.woodhead}@eedi.co.uk nancy@kitco.design
@@ 1.1 Contributions
  • We explore a variety of approaches to this task, including in-context learning, fine-tuning, and chain-of-thought prompting, together with rule- and sampling-based baselines. ***Our code is publicly available at https://github.com/umass-ml4ed/prompt_distractor_generation_NAACL
@@ 3.3 Experimental Setup
For all approaches except SB, we use a uniform format to represent the target MCQ. This format comprises a concatenation of 3 elements: the question stem, key, and its explanation. We use this structure since it encapsulates the most comprehensive information about the target MCQ. Furthermore, based on CoT, we instruct the LLM to first generate feedback message and then the distractor, which intends to simulate a reasoning pathway, providing a scaffold that guides the subsequent generation of plausible distractors. We use greedy decoding and a maximum output length of 350 tokens for distractor generation. Additional hyperparameters and model details are in Supplementary Material Section B. We also provide our prompts for CoT, RB, and kNN in Tables  LABEL:tab:CoT,  LABEL:tab:_RB, and  LABEL:tab:kNN respectively.
@@ 4 Conclusions and Future Work
In this paper, we explore automated distractor generation for math multiple-choice questions via large language models. We conduct experiments on a real-world math MCQ dataset and find that the in-context learning-based approach kNN, achieves the best performance when compared to other approaches such as fine-tuning, chain-of-thought prompting, and various baselines. We also conduct human evaluation and observe that LLMs are capable of generating mathematically valid distractors but are not fully aware of common errors or misconceptions among real students. Our initial exploration of this task opens up many avenues for future work. For example, we need to further refine the distribution-based metrics that predict the percentage of students who select each distractor. We also need to develop modified text encoding approaches that are closely aligned with errors or misconceptions among real students for in-context example selection. Furthermore, we aim to explore the generation of distractors, each of which corresponds to a specific error or misconception, in a planned and controllable way Wang et al. (2021); Zhang et al. (2023), possibly taking student knowledge state into account Liu et al. (2022), as well as the generation of high-quality feedback messages for each distractor Scarlatos et al. (2024). Finally, extending our work from multiple-choice questions to open-ended questions is important, since open-ended student responses contain much more detailed information on their errors Zhang et al. (2021, 2022); McNichols et al. (2023).
5 Acknowledgement
The authors thank Schmidt Futures and the NSF (under grants IIS-2118706 and IIS-2237676) for supporting this work.
@@ Appendix A Distribution ranking metric
Formally, we train an LLM-based model $\operatorname{r}_{\phi}(d_{1},d_{2},s,k,e_{k})\rightarrow\{d_{1},d_{2}\}$, where $\phi$ denotes the set of model parameters. We train this model by first constructing a dataset of all pairs of human-authored distractors for each MCQ and include both orders of each pair to avoid ordering bias, resulting in $N\times{3\choose 2}\times 2$ total pairs, where $N$ denotes the number of MCQs. Each pair is associated with a binary-valued label indicating whether $d_{1}$ or $d_{2}$ is selected by more students, which we can calculate from the student response records in our dataset. We then use this dataset to fine-tune an LLM in a text generation task, where the LLM receives the question and distractor information in its prompt and outputs its preference. We show our prompt for this task in Table LABEL:tab:prompt-ranking.
@@ Appendix C Human Evaluation Details
In this work, we obtained approval from the ethics review board for human evaluation. We show the evaluation instructions to human evaluators in Table  LABEL:tab:ins. We do not provide any compensation for human evaluators because their participation is entirely voluntary and we appreciate their contribution to this work.
@@ Appendix D Prompt Format
Prompt You are given the following math question along with the correct answer and explanation. Please use the following template to give 3 alternative incorrect answers to be used as multiple-choice options in a multiple-choice exam. Prior to the incorrect answer, provide feedback to be displayed to the student as an explanation of why that is not the correct answer.
@@
Prompt You are given the following math question along with the correct answer, explanation, and a list of errors. Please follow the template to first select 3 most likely errors for this question and use the selected errors to generate 3 alternative incorrect answers to be used as multiple-choice options in a multiple-choice exam. Prior to the incorrect answer, provide feedback to be displayed to the student as an explanation of why that is not the correct answer. If the list of errors is not given, generate 3 errors instead and do not contain any explanation in the 3 incorrect answer.
@@
Prompt Question: <in-context question>
@@ Appendix E Ranking Metric Examples
Prompt A teacher assigns the following math multiple choice question to a class of middle school students.
@@
Output Preferred Answer: A